I don't understand what I need to do here:
If log(x)=2.9 and log(y)=2.4, how do x and y compare?
The teacher gave us this problem, it is not from a book

Answer by nerdybill(7384)

If log(x)=2.9 and log(y)=2.4, how do x and y compare?
.
Solving for x:
log(x)=2.9
x = 10^2.9
x = 794.328
.
Solving for y:
log(y)=2.4
y = 10^2.4
y = 251.189
.
In other words, there is a difference of:
794.328 - 251.189 = 543.139
or percentage difference is:
251.189/794.328 * 100 = 31.6%